### v1

@@ 4 Transition from weak to energetic pulsars
Run $\mu_{*}\leavevmode\nobreak\ \left[\mathrm{m_{e}c^{2}R_{*}^{2}/e}\right]$ $\Omega_{*}\leavevmode\nobreak\ \left[\mathrm{c\leavevmode\nobreak\ rad/R_{*}}\right]$ $R_{s}\leavevmode\nobreak\ \left[\mathrm{R_{*}}\right]$ Injection $n_{\mathrm{inj}}$ $n_{0}$ $v_{\parallel}\leavevmode\nobreak\ \left[\mathrm{c}\right]$ $\sigma_{\mathrm{atm}}$ $k_{\mathrm{atm}}$ Resolution
@@
Table 2: Simulation parameters used for the second numerical experiment (Sect. 4), shown in Fig. 5.
@@ 5 Conclusions
Understanding neutron star magnetosphere dynamics is of utmost importance to reveal the locations of efficient particle acceleration and radiation generation. The connection between the rich pulsar observations and theoretical models is only possible through global ab initio simulations. In particular, due to the extreme nature of such astrophysical environments, kinetic models must capture self-consistent QED processes and general relativistic effects. This paper introduces a new general-relativistic module for the particle-in-cell code OSIRIS (Fonseca et al. 2002), intended to study plasma dynamics in intense electromagnetic and gravitational fields, capturing self-consistently GR effects in all components of the PIC algorithm (field solver, particle pusher, charge and current deposit scheme).
@@
We designed two numerical experiments to gain a deeper insight into the role of general relativity in activating the polar cap, focusing on the conditions for efficient particle acceleration. Therefore, we used the 4-norm of the current to probe under what circumstances the emission of a coherent radio beam along the magnetic axis is viable. In the first experiment, we injected plasma at $0.5\leavevmode\nobreak\ \left[\mathrm{c}\right]$ along the magnetic field lines, controlled by the local magnetization, to reproduce different quasi-stationary magnetospheric states. This setup allowed for a structured parametric scan over the plasma supply and stellar compactness. Without the general-relativistic effects, a low obliquity rotator would not be able to accelerate particles efficiently enough to potentiate the radio beam generation independently of the plasma supply. The general relativistic frame-dragging effect reduces the poloidal electric field, thus increasing the current amplitude. We show that this GR effect activates the polar cap of young and energetic neutron stars at any non-zero compactness, thus providing a robust mechanism for radio beam generation. For older neutron stars, it depends on both the compactness and plasma supply. Nevertheless, we show that GR could enable radio emission even for weak pulsars that drive weaker poloidal currents.
@@
In the second experiment (Sect. 4), we improve the accuracy of the previous model by resolving the polar cap gap dynamics using a heuristic ”on-the-spot” pair production model. Here, we allow pair production to populate only regions of efficient particle acceleration. The magnetosphere self-regulates to drive the required currents that sustain the magnetic field line torsion. By varying the pair production efficiency, i.e. through an increase in the magnetic field amplitude, we observe a clear transition between the weak and force-free pulsar solutions at fixed compactness (fiducial value of $R_{s}/R_{*}=0.3$). Once again, young and energetic neutron stars support strong poloidal currents, characteristic of a denser force-free magnetospheric solution, and possess active polar caps at any non-zero stellar compactnesses. Older pulsars, i.e. pulsars approaching their death line, transition to weak pulsar solutions and must rely on both the pair production efficiency (plasma supply) and stellar compactness for the presence of efficient particle acceleration and consequent radio beam. In Fig. 5, we demonstrate the existence of a minimum pair production efficiency for which the spacelike current develops in the bulk of the polar cap. The previous numerical experiment also captured this transient behaviour of the spacelike current via varying the plasma supply at fixed compactness (see Fig. 4). Also, simulations highlight the poloidal current-mediated interplay between the polar cap and outer gaps. Magnetospheres with such null surface gaps develop a volumetric return current, constricting the polar cap area and limiting the poloidal current in the open magnetic field line bundle. Consequently, equation (44) requires a lower poloidal electric field to activate the polar cap discharge, achieved by increasing the stellar compactness. Magnetospheres with pair production near the Y-point populate these exterior regions with plasma, relaxing the polar cap poloidal extension and pair production efficiency at the poles. This interconnection between outer and inner gaps may explain the intermittency of certain observed pulsars or even justify the existence of sub-pulse nullings.
@@
Future works with more accurate pair production models, i.e. heuristic models including non-zero photon mean free path and geodesic propagation, or even modelling QED processes more accurately with Monte Carlo techniques (Cruz et al. 2021), should explore this global outer-to-inner gap dynamics. Also, the variation of the upstream plasma supply from the stellar surface/atmosphere may affect the dynamics of the polar cap gap and its dependency on the driven magnetospheric current, requiring further investigation. Also, three-dimensional models of weak pulsar magnetospheres for low-obliquity rotators may unveil the importance of GR in the pulsar observational appearance.
Acknowledgements.
@@
This work is partially supported by the European Research Council (ERC-2015-AdG Grant 695088) and FCT (Foundation for Science and Technology, Portugal) under the project X-MASER No. 2022.02230.PTDC. RT is supported by FCT (Portugal) (Grant PD/BD/142971/2018) in the framework of the Advanced Program in Plasma Science and Engineering (APPLAuSE, FCT Grant PD/00505/2018). We acknowledge EuroHPC for granting access to LUMI (Large Unified Modern Infrastructure, Kajaani, Finland) within the EuroHPC-JU project EHPC-REG-2021R0038, where the simulations presented in this work were performed.
@@ Appendix B Charge deposition scheme
[×${\nabla}\times\left(\alpha\@vec{B}-\@vec{\beta}\times\@vec{E}\right)\right]=\@vec{\nabla}\cdot\left[\frac{\partial\@vec{E}}{\partial t}+4\pi\left(\alpha\@vec{j}-\rho\@vec{\beta}\right)\right],$ (66)
@@
$0=\frac{\partial\rho}{\partial t}+\@vec{\nabla}\cdot\left(\alpha\@vec{j}-\rho\@vec{\beta}\right)=\frac{\partial\rho}{\partial t}+\@vec{\nabla}\cdot\@vec{J}$ (67)
@@
Appendix C Mur open boundary
@@
The radiation condition impinging on a spherical shell placed far from the compact point source derived in (Sommerfeld 1949) reads
@@
describes only the outward propagating component of equation (80) and is hereafter designated as the Sommerfeld condition. At a finite distance from the source, it approximates to (Novak & Bonazzola 2004; Espinoza et al. 2014)

### v2

@@ 4 Transition from weak to energetic pulsars
Run $\mu_{*}~{}\left[\mathrm{m_{e}c^{2}R_{*}^{2}/e}\right]$ $\Omega_{*}~{}\left[\mathrm{c~{}rad/R_{*}}\right]$ $R_{s}~{}\left[\mathrm{R_{*}}\right]$ Injection $n_{\mathrm{inj}}$ $n_{0}$ $v_{\parallel}~{}\left[\mathrm{c}\right]$ $\sigma_{\mathrm{atm}}$ $k_{\mathrm{atm}}$ Resolution
@@
Table 2: Simulation parameters used for the second numerical experiment (Sect. 4), shown in Fig. 6. These simulations take on average 500k CPU core hours.
@@ 5 Conclusions
Understanding neutron star magnetosphere dynamics is of utmost importance to reveal the locations of efficient particle acceleration and radiation generation. The connection between the rich pulsar observations and theoretical models is only possible through global ab initio simulations. In particular, due to the extreme nature of such astrophysical environments, kinetic models must capture self-consistent QED processes and general relativistic effects. This paper introduces a new general-relativistic module for the particle-in-cell code OSIRIS (Fonseca et al., 2002), intended to study plasma dynamics in intense electromagnetic and gravitational fields, capturing self-consistently GR effects in all components of the PIC algorithm (field solver, particle pusher, charge and current deposit scheme).
@@
We designed two numerical experiments to gain a deeper insight into the role of general relativity in activating the polar cap, focusing on the conditions for efficient particle acceleration. Therefore, we used the 4-norm of the current to probe under what circumstances the emission of a coherent radio beam along the magnetic axis is viable. In the first experiment, we injected plasma at $0.5~{}\left[\mathrm{c}\right]$ along the magnetic field lines, controlled by the local magnetization, to reproduce different quasi-stationary magnetospheric states. This setup allowed for a structured parametric scan over the plasma supply and stellar compactness. Without the general-relativistic effects, a low obliquity rotator would not be able to accelerate particles efficiently enough to potentiate the radio beam generation independently of the plasma supply. The general relativistic frame-dragging effect reduces the poloidal electric field, thus increasing the current amplitude. We show that this GR effect activates the polar cap of young and energetic neutron stars at any non-zero compactness, thus providing a robust mechanism for radio beam generation. For older neutron stars, it depends on both the compactness and plasma supply. Nevertheless, we show that GR could enable radio emission even for weak pulsars that drive weaker poloidal currents. Therefore, weak pulsars may be observable in the radio spectrum beyond their expected death line, some of which can present themselves with narrower beams concerning force-free estimates. We note, however, that measuring such deviations is challenging due to other competing effects, such as the observer’s line of sight limitation or the misalignment angle $\chi$, requiring a full three-dimensional model extension. Previous kinetic studies of pulsar magnetospheres at arbitrary inclinations and low plasma injection rates demonstrated a similar behaviour of the magnetospheric solution (Kalapotharakos et al., 2018; Brambilla et al., 2018), with decreased Poynting flux and an outer gap beyond the null surface and around the separatrix region. Consequently, the volumetric field-aligned magnetospheric current and the torsion of the magnetic field lines decrease for these weak pulsar solutions. Following the conclusions from this work, we expect a reduction of the spacelike volume ($j_{\mu}j^{\mu}>0$ or $\alpha>1$) in the outflowing region of the polar cap (Bai and Spitkovsky, 2010; Philippov and Kramer, 2022), where pair production and the generation of the radio beam are expected. However, this effect should be more prominent in low-obliquity or very charge-starved pulsars, where this deviation from the force-free solution produces clear differences in the expected beam width. Our findings and conjecture seem to agree with observations of millisecond pulsars, which do not follow the scaling predicted from a canonical pulsar model for the inferred luminosity and angular radius of the generated radio beam (Kramer et al., 1998). The compactness of their magnetospheres and deviations from the force-free solution may explain such systematic tendencies.
@@
In the second experiment (Sect. 4), we improve the accuracy of the previous model by resolving the polar cap gap dynamics using a heuristic “on-the-spot” pair production model. Here, we allow pair production to populate only regions of efficient particle acceleration. The magnetosphere self-regulates to drive the required currents that sustain the magnetic field line torsion. By varying the pair production efficiency, i.e. through an increase in the magnetic field amplitude, we observe a clear transition between the weak and force-free pulsar solutions at fixed compactness (fiducial value of $R_{s}/R_{*}=0.3$). Once again, young and energetic neutron stars support strong poloidal currents, characteristic of a denser force-free magnetospheric solution, and possess active polar caps at any non-zero stellar compactnesses. Older pulsars, i.e. pulsars approaching their death line, transition to weak pulsar solutions and must rely on both the pair production efficiency (plasma supply) and stellar compactness for the presence of efficient particle acceleration and consequent radio beam. In Fig. 6, we demonstrate the existence of a minimum pair production efficiency for which the spacelike current develops in the bulk of the polar cap. The previous numerical experiment also captured this transient behaviour of the spacelike current via varying the plasma supply at fixed compactness (see Fig. 5). Also, simulations highlight the poloidal current-mediated interplay between the polar cap and outer gaps. Magnetospheres with such null surface gaps develop a volumetric return current, constricting the polar cap area and limiting the poloidal current in the open magnetic field line bundle. Consequently, equation (44) requires a lower poloidal electric field to activate the polar cap discharge, achieved by increasing the stellar compactness. Magnetospheres with pair production near the Y-point populate these exterior regions with plasma, relaxing the polar cap poloidal extension and pair production efficiency at the poles. This interconnection between outer and inner gaps may explain the intermittency of certain observed pulsars or even justify the existence of pulse nullings. In particular, weak pulsar solutions generate low-multiplicity plasma beamlets near the magnetic axis, compatible with the observed narrow-width dwarf pulses in Chen et al. (2023).
In this work, we have employed a novel reduced-domain approach to model the global magnetosphere of a neutron star. This reduced model allows for simulations with higher spatiotemporal resolution and better macro-particle statistics, at the expense of introducing an up-down equatorial symmetry. Consequently, by construction, this model does not capture the development of the kink instability in the equatorial current sheet (Cerutti et al., 2015; Philippov et al., 2015a). In that sense, we do not expect our method to model correctly the dissipation properties of thin current sheets, characterized by frequent magnetic reconnection events and plasmoid formation (Hakobyan et al., 2019, 2023; Schoeffler et al., 2023). However, we expect good agreement for weak pulsar solutions with no pair production near $R_{\mathrm{LC}}$, which possess a volumetric return-current and wide equatorial current sheet (with no significant dissipation/reconnection events occurring). Also, the location of the Y-point is consistent with similar full-domain particle-in-cell simulations (Chen et al., 2020; Guépin et al., 2020; Cruz et al., 2023; Hakobyan et al., 2023), indicating that this reduced model preserves the main elements of a neutron star magnetosphere in the aligned rotator configuration.
@@
Future works with more accurate pair production models, i.e. heuristic models including non-zero photon mean free path and geodesic propagation, or even modelling QED processes more accurately with Monte Carlo techniques (Cruz et al., 2021), should explore this global outer-to-inner gap dynamics. Also, the variation of the upstream plasma supply from the stellar surface/atmosphere may affect the dynamics of the polar cap gap and its dependency on the driven magnetospheric current, requiring further investigation. In particular, with the surface/atmospheric injection of electron-ion plasma instead of pair plasma, which affects the discharge dynamics on the return-current, modifying the plasma supply to the outer magnetosphere. Also, three-dimensional models of weak pulsar magnetospheres for low-obliquity rotators may unveil the importance of GR in the pulsar observational appearance.
Acknowledgements
@@
This work is partially supported by the European Research Council (ERC-2015-AdG Grant 695088) and FCT (Foundation for Science and Technology, Portugal) under the project X-MASER No. 2022.02230.PTDC. RT is supported by FCT (Portugal) (Grant PD/BD/142971/2018) in the framework of the Advanced Program in Plasma Science and Engineering (APPLAuSE, FCT Grant PD/00505/2018). We acknowledge EuroHPC for granting access to LUMI (Large Unified Modern Infrastructure, Kajaani, Finland) within the EuroHPC-JU project EHPC-REG-2021R0038, where the simulations presented in this work were performed. The authors thank the anonymous referee for the thoughtful and detailed comments that have significantly improved the manuscript.
@@ Appendix B Charge deposition scheme
[×$\boldsymbol{\nabla}\cdot\left[\boldsymbol{\nabla}\times\left(\alpha\boldsymbol{B}-\boldsymbol{\beta}\times\boldsymbol{E}\right)\right]=\boldsymbol{\nabla}\cdot\left[\frac{\partial\boldsymbol{E}}{\partial t}+4\pi\left(\alpha\boldsymbol{j}-\rho\boldsymbol{\beta}\right)\right],$ (57)
@@
$0=\frac{\partial\rho}{\partial t}+\boldsymbol{\nabla}\cdot\left(\alpha\boldsymbol{j}-\rho\boldsymbol{\beta}\right)=\frac{\partial\rho}{\partial t}+\boldsymbol{\nabla}\cdot\boldsymbol{J}$ (58)
@@
Appendix C Charge conservation
We demonstrate that the code conserves charge to machine precision independently of the selected coordinate system. To exemplify, we use the standard spherical coordinate system with a non-uniform grid spacing in both radial and meridional directions, i.e. uniform spacing in $\log r$ and $-\cos\theta$, and initialize particles with a thermal distribution of velocities in all directions, i.e. $\boldsymbol{u}_{\mathrm{thermal}}/c=(2.0,2.0,2.0)$, in the vicinity of a rotating neutron star with dipolar magnetic field $B_{*}=1000~{}[m_{e}c^{2}/eR_{*}]$ and $\Omega_{*}=0.1~{}[c~{}\mathrm{rad}/R_{*}]$, see upper-left panel in Fig. 7. Also, we selected a compactness parameter $R_{s}/R_{*}=0.5$ to show that the scheme is valid for significant curved spacetimes. Every iteration, we compute the deviation from the continuity equation given by
$\Delta_{\mathrm{continuity}}\equiv\frac{\partial\rho}{\partial t}+\boldsymbol{\nabla}\cdot\boldsymbol{J},$ (70)
$\Delta^{n+1/2}_{\mathrm{cont.}~{}i,j}\equiv\frac{\rho_{i,j}^{n+1}-\rho_{i,j}^{n}}{\Delta t}+\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{r},n+1/2}_{i,j}+\left(\boldsymbol{\nabla}\cdot\boldsymbol{J}\right)^{\hat{\theta},n+1/2}_{i,j},$ (71)
and evaluate the grid average value of that quantity, shown in the right panel of Fig. 7. Under the same conditions, but now using the logical coordinate system, i.e. using a uniform grid in $(\tilde{r},\tilde{\theta})=(\log r,-\cos\theta)$, yields the results in Fig. 8. Apart from small numerical fluctuations, the grid averaged deviation of the continuity equation stays close to machine precision over the test duration for both cases.
Figure 7: Charge conservation diagnostic for a randomized sample of electrons initialized with thermal velocities. This setup uses nonlinear grid spacings in both $r$ (logarithmic) and $\theta$ (equal area) to highlight that the proposed scheme works even when not using logical coordinate systems. Left panels show the initial and final states, the right panel displays the temporal evolution of the spatial averaged diagnostic.
Figure 8: Charge conservation diagnostic for a randomized sample of electrons initialized with thermal velocities. This setup uses uniform grid spacings in both $\tilde{r}$ and $\tilde{\theta}$, i.e. using the logical coordinate system. Left panels show the initial and final states, the right panel displays the temporal evolution of the spatial averaged diagnostic.
Appendix D Progressive filter
We explain in more detail the implementation of the generic progressive filter applied to the first and last radial domain cells. In particular, this progressive filter is used to filter the current density grid components, in this appendix designated $f_{i}^{k}$ with $i$ and $k$ being the cell index and the filtering order, respectively. Outside the simulation domain, the current density is unknown, thus invalidating high-order smoothing near the boundaries. Instead of not filtering the boundary cells or filling the guard cells with unphysical field values, we compute the exterior current component values such that only the first and last cells remain unfiltered.
Within the OSIRIS framework, smoothing any field component to order $n$ requires a single-passage kernel with $12n+$ 1 entries. For example, the single-passage first-, second-, and third-order kernels read
$\mathcal{K}^{1}=\left[K^{1}_{1}\;\;K^{1}_{2}\;\;K^{1}_{3}\right],$ (72)
$\mathcal{K}^{2}=\left[K^{2}_{1}\;\;K^{2}_{2}\;\;K^{2}_{3}\;\;K^{2}_{4}\;\;K^{2}_{5}\right],$ (73)
$\mathcal{K}^{3}=\left[K^{3}_{1}\;\;K^{3}_{2}\;\;K^{3}_{3}\;\;K^{3}_{4}\;\;K^{3}_{5}\;\;K^{3}_{6}\;\;K^{3}_{7}\right],$ (74)
with $K$ being the weight given to each neighbour. The third-order filtered value of the first four cells yields
$\color[rgb]{0,0,0}f^{3}_{1}=K^{3}_{1}\color[rgb]{1,0,0}f^{0}_{-2}\color[rgb]{0,0,0}+K^{3}_{2}\color[rgb]{1,0,0}f^{0}_{-1}\color[rgb]{0,0,0}+K^{3}_{3}\color[rgb]{1,0,0}f^{0}_{0}\color[rgb]{0,0,0}+K^{3}_{4}f^{0}_{1}+K^{3}_{5}f^{0}_{2}+K^{3}_{6}f^{0}_{3}+K^{3}_{7}f^{0}_{4},$ (75)
$\color[rgb]{0,0,0}f^{3}_{2}=K^{3}_{1}\color[rgb]{1,0,0}f^{0}_{-1}\color[rgb]{0,0,0}+K^{3}_{2}\color[rgb]{1,0,0}f^{0}_{0}\color[rgb]{0,0,0}+K^{3}_{3}f^{0}_{1}+K^{3}_{4}f^{0}_{2}+K^{3}_{5}f^{0}_{3}+K^{3}_{6}f^{0}_{4}+K^{3}_{7}f^{0}_{5},$ (76)
$\color[rgb]{0,0,0}f^{3}_{3}=K^{3}_{1}\color[rgb]{1,0,0}f^{0}_{0}\color[rgb]{0,0,0}+K^{3}_{2}f^{0}_{1}+K^{3}_{3}f^{0}_{2}+K^{3}_{4}f^{0}_{3}+K^{3}_{5}f^{0}_{4}+K^{3}_{6}f^{0}_{5}+K^{3}_{7}f^{0}_{6},$ (77)
$\color[rgb]{0,0,0}f^{3}_{4}=K^{3}_{1}f^{0}_{1}+K^{3}_{2}f^{0}_{2}+K^{3}_{3}f^{0}_{3}+K^{3}_{4}f^{0}_{4}+K^{3}_{5}f^{0}_{5}+K^{3}_{6}f^{0}_{6}+K^{3}_{7}f^{0}_{7},$ (78)
where the terms in red are the unknown field values inside the guard cells. Equations (75)-(78) highlight that we can filter the fourth cell using the third-order single passage filter in eq. (74), the third cell using the second-order single passage filter in eq. (73), the second cell using the first-order single-passage filter in eq. (72), and the first cell remains unfiltered, i.e.
$f^{3}_{3}=f_{3}^{2}=K^{2}_{1}f^{0}_{1}+K^{2}_{2}f^{0}_{2}+K^{2}_{3}f^{0}_{3}+K^{2}_{4}f^{0}_{4}+K^{2}_{5}f^{0}_{5},$ (79)
$f^{3}_{2}=f_{2}^{1}=K^{1}_{1}f^{0}_{1}+K^{1}_{2}f^{0}_{2}+K^{1}_{3}f^{0}_{3},$ (80)
$f^{3}_{1}=f_{1}^{0}.$ (81)
Inserting eqs. (75)-(77) into eqs. (79)-(81) determines the three unknown field components defined in the guard cells, i.e.
$f_{0}^{0}=\frac{1}{K_{1}^{3}}\left[\left(K_{1}^{2}-K_{2}^{3}\right)f^{0}_{1}+\left(K_{2}^{2}-K_{3}^{3}\right)f^{0}_{2}+\right.$
$+\left.\left(K_{3}^{2}-K_{4}^{3}\right)f^{0}_{3}+\left(K_{4}^{2}-K_{5}^{3}\right)f^{0}_{4}+\left(K_{5}^{2}-K_{6}^{3}\right)f^{0}_{5}-K_{7}^{3}f^{0}_{6}\right],$ (82)
$f_{-1}^{0}=\frac{1}{K_{1}^{3}}\left[-K_{2}^{3}f^{0}_{0}+\left(K_{1}^{1}-K_{3}^{3}\right)f^{0}_{1}+\left(K_{2}^{1}-K_{4}^{3}\right)f^{0}_{2}+\right.$
$+\left.\left(K_{3}^{1}-K_{5}^{3}\right)f^{0}_{3}-K_{6}^{3}f^{0}_{4}-K_{7}^{3}f^{0}_{5}\right],$ (83)
$f_{-2}^{0}=\frac{1}{K_{1}^{3}}\left[-K_{2}^{3}f_{-1}^{0}-K_{3}^{3}f_{0}^{0}+\left(1-K_{4}^{3}\right)f_{1}^{0}-K_{5}^{3}f_{2}^{0}-\right.$
$-\left.K_{6}^{3}f_{3}^{0}-K_{7}^{3}f_{4}^{0}\right].$ (84)
All other field values are filtered using the third-order single passage kernel.
Appendix E Mur open boundary
@@
The radiation condition impinging on a spherical shell placed far from the compact point source derived in (Sommerfeld, 1949) reads
@@
describes only the outward propagating component of equation (86) and is hereafter designated as the Sommerfeld condition. At a finite distance from the source, it approximates to (Novak and Bonazzola, 2004; Espinoza et al., 2014)